OpenStudy (anonymous):

a square indoor play space has an are of 121ft^2. what is the perimeter of the play space? please explain

OpenStudy (anonymous):

area of square = side^2 therefore side of play graound=(121)^0.5 = 11 so perimeter = 11*4=44 since all sides of the square are equale

OpenStudy (anonymous):

what does this * mean

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

wich number is the exponent

OpenStudy (anonymous):

121^0.5 means squre root of 121

OpenStudy (anonymous):

ok thanks